multiply. (7-9x^6)^2
Answer by HyperBrain(694)

Hmmm...
Look here. IF I have a binomial a-b and I square it,

(a-b)2=(a-b)(a-b)

(a-b)2=a(a-b)-b(a-b)

(a-b)2=(a2-ab)-(ab-b2)

(a-b)2=a2-ab-ab+b2

(a-b)2=a2-2ab++b2

Applying here,

(7-9x6)2=72-2(7)(9x6)+(9x6)2

(7-9x6)2=49-(2*7*9)(x6)+(92)(x6)2

(7-9x6)2=49-126x6+81x12

Power up,
HyperBrain!